OpenStudy (anonymous):

A ladder 42 feet in length rests against a vertical wall of 36 feet in height. What is the measure of the angle of elevation, θ. (Round angle measure to the nearest degree.) 31° 59° 41°

OpenStudy (anonymous):



OpenStudy (anonymous):

Use: \[\sin(\theta) = \frac{Perpendicular}{Hypotenuse} = \frac{36}{42}\] Find \(\theta\) here..

OpenStudy (anonymous):

58.9 or 59?

OpenStudy (anonymous):

Sorry net got disconnected.. Yep 58.99 or 59 is right..

OpenStudy (anonymous):

Thanks for helping me. I totally understand. That has been happening a lot for me.